f(x)= x^2+3x+b (x>1,X=1)


         2ax+3       (x<1)


f(x) is continous and differenstiable then find value of a and b

Grade:12

1 Answers

Ramesh V
14 years ago

at x=1 f(1) = 1+3+b=b+4

at x=1-h whr h-->0 , f(1-h)=2a(1-h)+3=2a+3

at x=1+h whr h-->0 , f(1+h)=(1+h)^2+3(1+h)+b=1+3+b

as all 3 eqns shld be equal for function to be continuous, it means : 2a+3 = b+4

f(x) is differentiable

f'(x) = 2x+3 at x=1 its f'(1)=5

f'(x+h)= [ f(x+h)-f(x)] / h = [ f(1+h)-f(1)] / h = [ (1+h)^2+3(1+h)+b - b-4 ] / h = h+5 = 5

f'(x-h)=[ f(x-h)-f(x) ]/ -h = [ f(1-h)-f(1)] / -h = [ 2a -2ah +3 - b-4 ] / h

to cancel on both numerator and denominator, (b+1-2a) shld be 0

 = 2a = 5

a=5/2 .. on putting above we have

b=4

--

regards

Ramesh